# ModernGBERT: German-only 1B Encoder Model Trained from Scratch

### v1

@@ 5.1 Intermediate Model Evaluation
These saturation patterns, including per-task trends and overall performance plateaus, are consistent with findings by Pfister et al. (2024) for decoder models, and by Antoun et al. (2024) for their ModernBERT variant ModernCamemBERT (with 136M parameters) trained on French. Our results confirm that although small ModernBERT models saturate quickly, larger models benefit from additional data. Extrapolating the observed scaling behavior between ModernGBERT 134M and 1B, we hypothesize that training a larger, 7B-sized encoder could make further use of an extensive monolingual datasets, improving performance beyond ModernGBERT 1B.
5.2 Final Model Evaluation
Model | Avg | Class. | NER | PAWSX | QA |
GBERTBase | 0.718 | 0.723 | 0.786 | 0.561 | 0.803 |
GBERTLarge | 0.768 | 0.785 | 0.799 | 0.654 | 0.832 |
GeBERTaBase | 0.716 | 0.715 | 0.778 | 0.559 | 0.813 |
GeBERTaLarge | 0.749 | 0.743 | 0.791 | 0.619 | 0.844 |
GeBERTaXLarge | 0.767 | 0.770 | 0.807 | 0.643 | 0.848 |
XLM-RoBERTaBase | 0.689 | 0.693 | 0.754 | 0.505 | 0.802 |
XLM-RoBERTaLarge | 0.730 | 0.714 | 0.787 | 0.583 | 0.837 |
XLM-RoBERTaXLarge | 0.758 | 0.750 | 0.802 | 0.656 | 0.822 |
LLäMmlein 120M | 0.676 | 0.702 | 0.712 | 0.477 | 0.812 |
LLäMmlein2Vec 120M | 0.684 | 0.703 | 0.741 | 0.472 | 0.819 |
LLäMmlein 1B | 0.733 | 0.781 | 0.773 | 0.548 | 0.828 |
LLäMmlein2Vec 1B | 0.762 | 0.776 | 0.812 | 0.615 | 0.843 |
LLäMmlein 7B | 0.747 | 0.810 | 0.805 | 0.524 | 0.851 |
LLäMmlein2Vec 7B | 0.787 | 0.799 | 0.838 | 0.670 | 0.842 |
ModernGBERT 134M◇ | 0.730 | 0.716 | 0.782 | 0.589 | 0.833 |
ModernGBERT 134M | 0.749 | 0.735 | 0.805 | 0.612 | 0.836 |
ModernGBERT 1B◇ | 0.800 | 0.806 | 0.839 | 0.681 | 0.874 |
ModernGBERT 1B | 0.808 | 0.812 | 0.845 | 0.699 | 0.876 |
@@ 5.2 Final Model Evaluation
We evaluate all final models on the full SuperGLEBer benchmark. Table 2 averages the scores per task type, while Table 7 provides more fine-grained scores. We compare our models to established encoders: GBERT (Chan et al., 2020), GeBERTa (Dada et al., 2023), and XLM-RoBERTa (Conneau et al., 2020; Goyal et al., 2021).
@@
Our ModernGBERT consistently outperforms comparable and larger models. The 134M variant achieves an average score of 0.749, surpassing all similar-sized baselines, including GBERTBase (0.718), XLM-RoBERTaBase (0.689), GeBERTaBase (0.716), and even XLM-RoBERTaLarge (0.730), as well as LLäMmlein 1B (0.733). The ModernGBERT 1B variant achieves a new state-of-the-art average score across the entire SuperGLEBer of 0.808, outperforming GBERTLarge (0.768) by 4% and beating the seven times larger LLäMmlein2Vec 7B (0.787). It leads in three of four evaluation categories, including classification (0.812), NER (0.845), and QA (0.876). Only on sentence similarity (0.699), our seven times larger LLäMmlein2Vec 7B achieves better results. ModernGBERT scales well, with performance improving for larger model sizes, again suggesting that scaling ModernBERT-style encoders can leverage large monolingual corpora effectively. In the SuperGLEBer setting, adding context extension improved ModernGBERT’s average by 1.9% for the 134M model (from 0.730 to 0.749) and by 0.8% for the 1B variant (from 0.800 to 0.808). No large improvements were to be expected from our context extension, as SuperGLEBer tasks do not make use of long contexts.
@@
Adaptation via LLM2Vec yields consistent gains across models. Thus, our first LLM2Vec tuning (analogous to ext1, Section 2.2) showed the most prominent positive effect, while the second finetune using the ext2 datasets showed only marginal increase, or even a decrease in performance. The same holds for a mixture of the two LLM2Vec adapters (ext1+2). The LLäMmlein2Vec 7B achieves the strongest results among the LLM2Vec models (0.787). Conversion of LLäMmlein 120M, 1B, 7B improved the average score by 0.8%, 2.9%, and 4.0% respectively. This effect is especially pronounced in PAWSX, with scores increasing by up to 14.6% for LLäMmlein 7B and 6.7% for LLäMmlein 1B.
@@
Comparing the LLäMmlein2Vec with the ModernGBERT family, we find that on similarly sized models, ModernGBERT always outperforms the transformed decoders by a large margin. Only the much larger LLäMmlein2Vec 7B approaches the performance of ModernGBERT 1B.
Model | Avg | Clustering | Reranking | Retrieval |
GBERTBase | 0.360 | 0.274 | 0.118 | 0.226 |
GBERTBase † | 0.500 | 0.318 | 0.374 | 0.461 |
GBERTLarge | 0.412 | 0.336 | 0.206 | 0.297 |
GBERTLarge † | 0.521 | 0.334 | 0.389 | 0.493 |
XLM-RoBERTaBase | 0.248 | 0.173 | 0.024 | 0.008 |
XLM-RoBERTaBase† | 0.403 | 0.247 | 0.247 | 0.299 |
XLM-RoBERTaLarge | 0.264 | 0.172 | 0.048 | 0.026 |
XLM-RoBERTaLarge† | 0.460 | 0.259 | 0.343 | 0.416 |
XLM-RoBERTaXLarge | 0.301 | 0.225 | 0.090 | 0.142 |
XLM-RoBERTaXLarge† | 0.479 | 0.342 | 0.362 | 0.407 |
LLäMmlein2Vec 120M | 0.315 | 0.261 | 0.139 | 0.224 |
LLäMmlein2Vec 120M† | 0.471 | 0.308 | 0.325 | 0.425 |
LLäMmlein2Vec 1B | 0.399 | 0.308 | 0.183 | 0.276 |
LLäMmlein2Vec 1B† | 0.540 | 0.343 | 0.433 | 0.511 |
LLäMmlein2Vec 7B | 0.376 | 0.249 | 0.169 | 0.266 |
LLäMmlein2Vec 7B† | 0.557 | 0.339 | 0.477 | 0.522 |
ModernGBERT 134M | 0.383 | 0.293 | 0.139 | 0.241 |
ModernGBERT 134M◇† | 0.485 | 0.303 | 0.364 | 0.432 |
ModernGBERT 134M | 0.376 | 0.296 | 0.120 | 0.213 |
ModernGBERT 134M† | 0.501 | 0.312 | 0.404 | 0.446 |
ModernGBERT 1B | 0.374 | 0.318 | 0.097 | 0.199 |
ModernGBERT 1B◇† | 0.549 | 0.339 | 0.463 | 0.511 |
ModernGBERT 1B | 0.366 | 0.307 | 0.088 | 0.191 |
ModernGBERT 1B† | 0.551 | 0.338 | 0.459 | 0.512 |
@@
We evaluate models on the MTEB benchmark, which covers six task categories: classification, pair classification, clustering, reranking, retrieval, and short text similarity (STS) tasks. While Table 3 summarizes the outcomes, all results are presented in Table 9. In general, supervised fine-tuning on mMARCO yields consistent improvements across all model types. While classification performance sometimes declines, substantial gains can be observed in other areas: 25% on average for reranking, 26% for retrieval and 25% for STS.
The best overall average performance is achieved by the fine-tuned LLäMmlein2Vec 7B (0.557), closely followed by the fine-tuned ModernGBERT 1B (0.551), despite the latter being significantly smaller. LLäMmlein2Vec models generally show strong performance after fine-tuning, particularly when trained with the extension dataset of the first phase (ext1). Using the second extension phase (ext2) or combining both adapters into the base model (ext1+2) harms the performance. Interestingly, the latter shows the largest fine-tuning gains among the three variants.
The ModernGBERT models perform competitively to similarly sized models. Before fine-tuning, ModernGBERT 1B (avg. 0.366) already outperforms most encoder-only models, such as GeBERTaXLarge (0.325) or XLM-RoBERTaXLarge (0.301), but not GBERTLarge (0.412). However, after fine-tuning, it demonstrates clear superiority among native encoder-only models by at least 3% on the average score. As with our observations on the SuperGLEBer benchmark, ModernGBERT’s context extension did not show significant improvements here.
Comparing ModernGBERT and LLäMmlein2Vec, we find that before fine-tuning, the LLäMmlein2Vec 1B and 7B models produce better representations than ModernGBERT 1B. However, after fine-tuning, ModernGBERT 1B surpasses the 1B variant of LLäMmlein2Vec on average and closely aligns with the larger 7B model.

### v2

@@ 5.1. Intermediate Model Evaluation
These saturation patterns, including per-task trends and overall performance plateaus, are consistent with findings by Pfister et al. (2024) for decoder models, and by Antoun et al. (2024) for the French ModernBERT variant ModernCamemBERT (136M parameters). Our results confirm that while small ModernBERT models saturate quickly, larger models continue to benefit from additional data. Extrapolating from this observed scaling behaviour between ModernGBERT 134M and 1B, we hypothesize that a larger 7B encoder could leverage extensive monolingual corpora to achieve performance beyond ModernGBERT 1B.
LLäMmlein2Vec Conversion
To remain as consistent as possible with the ModernGBERT training procedure, we trained the LLäMmlein2Vec conversions on the same post-training datasets. To assess whether training on the full dataset was necessary, we evaluated multiple checkpoints throughout the entire training process on the six SuperGLEBer tasks. In the bottom half of Figure˜2 we illustrate the average SuperGLEBer performance alongside the of the standard decoder models (blue), as well as the performance of decoders with a bidirectional attention mask (purple). Interestingly, for encoder-typical tasks such as sentence similarity (PAWSX), the performance of both the 120M, 1B and 7B LLäMmlein models increased even without explicit bidirectional continued pre-training, showing gains from bidirectional fine-tuning alone (120M: 0.477 0.516; 1B: 0.548 0.580; 7B: 0.524 0.660).
Regarding the increases in average performance, the scores remained largely consistent over time, without showing significant improvements. This suggests that training could have been stopped earlier, thus reducing the overall training time, emphasizing that converting a decoder is substantially faster than training a model from scratch.
5.2. Final Model Evaluation
| Model | Size | Avg. | Class. | NER | PAWSX | QA |
| GBERTBase | 111M | 0.718 | 0.723 | 0.786 | 0.561 | 0.803 |
| GBERTLarge | 337M | 0.768 | 0.785 | 0.799 | 0.654 | 0.832 |
| GeBERTaBase | 139M | 0.716 | 0.715 | 0.778 | 0.559 | 0.813 |
| GeBERTaLarge | 406M | 0.749 | 0.743 | 0.791 | 0.619 | 0.844 |
| GeBERTaXLarge | 887M | 0.767 | 0.770 | 0.807 | 0.643 | 0.848 |
| XLM-RoBERTaBase | 279M | 0.689 | 0.693 | 0.754 | 0.505 | 0.802 |
| XLM-RoBERTaLarge | 561M | 0.730 | 0.714 | 0.787 | 0.583 | 0.837 |
| XLM-RoBERTaXLarge | 3.48B | 0.758 | 0.750 | 0.802 | 0.656 | 0.822 |
| mmBERT-small | 142M | 0.748 | 0.759 | 0.800 | 0.606 | 0.828 |
| mmBERT-base | 309M | 0.758 | 0.757 | 0.804 | 0.622 | 0.849 |
| EuroBERT-210m | 212M | 0.706 | 0.659 | 0.698 | 0.607 | 0.859 |
| EuroBERT-610m | 609M | 0.721 | 0.633 | 0.750 | 0.633 | 0.869 |
| EuroBERT-2b | 2.11B | 0.731 | 0.633 | 0.764 | 0.655 | 0.873 |
| LLäMmlein | 120M | 0.676 | 0.702 | 0.712 | 0.477 | 0.812 |
| LLäMmlein2Vec (ext1) | 120M | 0.684 | 0.703 | 0.741 | 0.472 | 0.819 |
| LLäMmlein | 1B | 0.733 | 0.781 | 0.773 | 0.548 | 0.828 |
| LLäMmlein2Vec (ext1) | 1B | 0.762 | 0.776 | 0.812 | 0.615 | 0.843 |
| LLäMmlein | 7B | 0.747 | 0.810 | 0.805 | 0.524 | 0.851 |
| LLäMmlein2Vec (ext1) | 7B | 0.787 | 0.799 | 0.838 | 0.670 | 0.842 |
| ModernGBERT | 134M | 0.730 | 0.716 | 0.782 | 0.589 | 0.833 |
| ModernGBERT (ext1+2) | 134M | 0.749 | 0.735 | 0.805 | 0.612 | 0.836 |
| ModernGBERT | 1B | 0.800 | 0.806 | 0.839 | 0.681 | 0.874 |
| ModernGBERT (ext1+2) | 1B | 0.808 | 0.812 | 0.845 | 0.699 | 0.876 |
@@ 5.2. Final Model Evaluation
We evaluate all final models on the full SuperGLEBer benchmark. Table˜2 reports average performances across a subset of our models. In particular, we compare our models to established encoders: GBERT (Chan et al., 2020), GeBERTa (Dada et al., 2023), XLM-RoBERTa (Conneau et al., 2020; Goyal et al., 2021), mmBERT Marone et al. (2025) and EuroBERT (Boizard et al., 2025).
@@
Our ModernGBERT consistently outperforms comparable and larger models. The 134M variant achieves an average score of 0.749, surpassing all similar-sized baselines, as well as LLäMmlein 1B (0.733), EuroBERT-2b (0.731) and LLäMmlein 7B (0.747). The ModernGBERT 1B variant achieves a new state-of-the-art average score across the entire SuperGLEBer of 0.808, outperforming GBERTLarge (0.768) by 4% and beating the seven times larger LLäMmlein2Vec 7B (0.787). It leads in three of four evaluation categories, including classification (0.812), NER (0.845), and QA (0.876). Only on sentence similarity (0.699), LLäMmlein2Vec 7B achieves better results. ModernGBERT scales well, with performance improving for larger model sizes, again suggesting that scaling ModernBERT-style encoders can leverage large monolingual corpora effectively. In the SuperGLEBer setting, adding context extension improved ModernGBERT’s average by 1.9% for the 134M model (from 0.730 to 0.749) and by 0.8% for the 1B variant (from 0.800 to 0.808). However, large improvements were not expected, as SuperGLEBer tasks do not make use of long contexts.
@@
Adaptation via LLM2Vec yields consistent gains across models. Our first LLM2Vec tuning (on ext1, Section˜2.2) showed the most prominent positive effect, while the second fine-tune using the ext2 datasets showed only marginal increase, or even a decrease in performance. The same holds for a mixture of the two LLM2Vec adapters (ext1+2). Therefore, we report only ext1 results in this and all subsequent tables for simplicity. The LLäMmlein2Vec 7B achieves the strongest results among the LLM2Vec models (0.787). Conversion of LLäMmlein 120M, 1B, 7B improved the average score by 0.8%, 2.9%, and 4.0% respectively. This effect is especially pronounced in PAWSX, with scores increasing by up to 14.6% for LLäMmlein 7B and 6.7% for LLäMmlein 1B.
@@
Comparing the LLäMmlein2Vec with the ModernGBERT family, we find that on similarly sized models, ModernGBERT always outperforms the transformed decoders by a large margin. Only the much larger LLäMmlein2Vec 7B approaches the performance of ModernGBERT 1B. This systematic comparison using identical datasets provides the first comprehensive analysis of MLM vs. LLM2Vec for encoder development.
| Model | Size | Avg. | Clstr | ReRnk | Retr |
| GBERTBase | 111M | 0.360 | 0.274 | 0.118 | 0.226 |
| GBERTBase f | 111M | 0.500 | 0.318 | 0.374 | 0.461 |
| GBERTLarge | 337M | 0.412 | 0.336 | 0.206 | 0.297 |
| GBERTLarge f | 337M | 0.521 | 0.334 | 0.389 | 0.493 |
| GeBERTaBase | 139M | 0.382 | 0.312 | 0.174 | 0.213 |
| GeBERTaBase f | 139M | 0.493 | 0.318 | 0.374 | 0.430 |
| GeBERTaLarge | 406M | 0.397 | 0.287 | 0.223 | 0.274 |
| GeBERTaLarge f | 406M | 0.494 | 0.311 | 0.374 | 0.432 |
| GeBERTaXLarge | 887M | 0.325 | 0.278 | 0.108 | 0.058 |
| GeBERTaXLarge f | 887M | 0.521 | 0.323 | 0.414 | 0.462 |
| XLM-RoBERTaBase | 279M | 0.248 | 0.173 | 0.024 | 0.008 |
| XLM-RoBERTaBase f | 279M | 0.403 | 0.247 | 0.247 | 0.299 |
| XLM-RoBERTaLarge | 561M | 0.264 | 0.172 | 0.048 | 0.026 |
| XLM-RoBERTaLarge f | 561M | 0.460 | 0.259 | 0.343 | 0.416 |
| XLM-RoBERTaXLarge | 3.48B | 0.301 | 0.225 | 0.090 | 0.142 |
| XLM-RoBERTaXLarge f | 3.48B | 0.479 | 0.342 | 0.362 | 0.407 |
| mmBERT-small | 142M | 0.289 | 0.163 | 0.091 | 0.075 |
| mmBERT-small f | 142M | 0.430 | 0.167 | 0.390 | 0.359 |
| mmBERT-base | 309M | 0.318 | 0.237 | 0.092 | 0.071 |
| mmBERT-base f | 309M | 0.475 | 0.214 | 0.428 | 0.406 |
| EuroBERT-210m | 212M | 0.293 | 0.240 | 0.070 | 0.118 |
| EuroBERT-210m f | 212M | 0.431 | 0.189 | 0.404 | 0.371 |
| EuroBERT-610m | 609M | 0.299 | 0.232 | 0.093 | 0.134 |
| EuroBERT-610m f | 609M | 0.419 | 0.288 | 0.263 | 0.352 |
| EuroBERT-2b | 2.11B | 0.230 | 0.186 | 0.060 | 0.065 |
| EuroBERT-2b f | 2.11B | 0.452 | 0.191 | 0.460 | 0.420 |
| LLäMmlein2Vec (ext1) | 120M | 0.315 | 0.261 | 0.139 | 0.224 |
| LLäMmlein2Vec (ext1)f | 120M | 0.471 | 0.308 | 0.325 | 0.425 |
| LLäMmlein2Vec (ext1) | 1B | 0.399 | 0.308 | 0.183 | 0.276 |
| LLäMmlein2Vec (ext1)f | 1B | 0.540 | 0.343 | 0.433 | 0.511 |
| LLäMmlein2Vec (ext1) | 7B | 0.376 | 0.249 | 0.169 | 0.266 |
| LLäMmlein2Vec (ext1)f | 7B | 0.557 | 0.339 | 0.477 | 0.522 |
| ModernGBERT | 134M | 0.383 | 0.293 | 0.139 | 0.241 |
| ModernGBERT f | 134M | 0.485 | 0.303 | 0.364 | 0.432 |
| ModernGBERT (ext1+2) | 134M | 0.376 | 0.296 | 0.120 | 0.213 |
| ModernGBERT (ext1+2) f | 134M | 0.501 | 0.312 | 0.404 | 0.446 |
| ModernGBERT | 1B | 0.374 | 0.318 | 0.097 | 0.199 |
| ModernGBERT f | 1B | 0.549 | 0.339 | 0.463 | 0.511 |
| ModernGBERT (ext1+2) | 1B | 0.366 | 0.307 | 0.088 | 0.191 |
| ModernGBERT (ext1+2) f | 1B | 0.551 | 0.338 | 0.459 | 0.512 |
@@
We evaluate models on the MTEB benchmark, which covers six task categories: classification, pair classification, clustering, reranking, retrieval, and short text similarity (STS) tasks. A summary of results is visualized in Table˜3. In general, supervised fine-tuning on mMARCO yields consistent improvements across all model types. While classification performance sometimes declines, substantial gains can be observed in other areas: 25% on average for reranking, 26% for retrieval and 25% for STS.
The best overall average performance is achieved by the fine-tuned LLäMmlein2Vec 7B (0.557), closely followed by the fine-tuned ModernGBERT 1B (0.551), despite the latter being significantly smaller. LLäMmlein2Vec models generally show strong performance after fine-tuning, particularly when trained with the extension dataset of the first phase (ext1). Using the second extension phase (ext2) or combining both adapters into the base model (ext1+2) again harms the performance. Interestingly, the latter shows the largest fine-tuning gains among the three variants. The ModernGBERT models perform competitively to similarly sized models. Before fine-tuning, ModernGBERT 1B (avg. 0.366) already outperforms most encoder-only models, such as GeBERTaXLarge (0.325), XLM-RoBERTaXLarge (0.301), and EuroBERT-2b (0.230) but not GBERTLarge (0.412). However, after fine-tuning, it demonstrates clear superiority among native encoder-only models by at least 3% on the average score. As with our observations on the SuperGLEBer benchmark, ModernGBERT’s context extension did not show significant improvements here. Comparing ModernGBERT and LLäMmlein2Vec, we find that before fine-tuning, the LLäMmlein2Vec 1B and 7B models produce better representations than ModernGBERT 1B. However, after fine-tuning, ModernGBERT 1B surpasses the 1B variant of LLäMmlein2Vec on average and closely aligns with the larger 7B model.